# Long-range Ising model for regional-scale seismic risk analysis

### v1

@@ 4 Parametric study of the Ising model
This section investigates the relationships between the Ising model parameters and typical performance quantities, such as failure probability and the correlation coefficient between damage states of structures, using simple proof-of-concept examples.
@@ 4.1 Interpretations of the Ising model parameters
    $\boldsymbol{h}$: We interpret $\boldsymbol{h}$ as a risk field acting on each structure within a region. According to Eq. (9), a structure with a positive $h_{i}$ is more probable to fail, i.e., to have a spin of $x_{i}=+1$, with a larger $h_{i}$ signifying a greater tendency towards failure. Conversely, a negative risk field $h_{i}$ indicates a tendency to remain safe. Given that civil structures are engineered to be reliable, they should inherently possess a negative field $-\mathcal{I}$, with $\mathcal{I}>0$. A hazard imposes an external positive field $h_{H}>0$ on each structure, and the resulting risk field is a balance between the hazard and the inertia towards remaining safe, i.e., $h_{i}=h_{H}-\mathcal{I}$. Here, both $h_{H}$ and $\mathcal{I}$ vary across structures.
@@
    $\boldsymbol{J}$: This represents the pairwise interactions, reflecting the tendency of two structures to be in the same damage state. The term interaction encompasses all direct and indirect effects stemming from physical and statistical properties that determine how the damage state of one structure can affect the state of another. From Eq. (9), it is observed that if $J_{ij}$ is positive, the probability of two structures being in the same state is higher than them being in different states, with a larger $J_{ij}$ increasing this probability. On the other hand, a negative $J_{ij}$ favors the damage states being different.
The specific choice of terminology, such as risk field and pairwise interaction, can vary, but the fundamental implications of $\boldsymbol{h}$ and $\boldsymbol{J}$ remain invariant. Recall from Section 3 that the construction of the Ising model relies solely on the first- and second-order cross-moments. Therefore, it seems valid to claim that the Ising model cannot offer anything beyond the mean and the covariance matrix. However, it is important to acknowledge that the principle of maximum entropy introduces a crucial additional layer of information. Therefore, the Ising model encompasses more than just the mean and the covariance matrix.
@@ 4.2 Dependency between the Ising model parameters and the mean and correlation coefficient
Consider two binary random variables $\boldsymbol{X}=[X_{1},X_{2}]^{\intercal}\in\{-1\text{(safe)},1\text{(failure)}\}^{2}$ representing the damage states of two buildings. The risk field and pairwise interaction are given as $h_{1}=h_{2}=h$ and $J_{12}=J$, respectively. With only 2 variables and thus $2^{2}$ system configurations, the exact mean and correlation coefficient for given $h$ and $J$ values can be determined using Eqs. (5)-(9). The mean value $\mathbb{E}\left[X_{i}\right]$ is directly linked to the failure probability $P_{f}$ by $P_{f}=(\mathbb{E}\left[X_{i}\right]+1)/2$. Note that, due to parameter settings of the current demonstration, the failure probability $P_{f}$ for each building is the same.
The mean value for the damage state of a building is affected both by the risk field acting on the site and the pairwise interaction from the other building. Figure 0(a) depicts the mean damage state of a building with varying $h$ and $J$, and Figure 0(b) illustrates the relationship between the mean value and the risk field $h$ under four different $J$ values. The mean damage state of a building becomes larger as the risk field $h$ increases, reflecting that a stronger risk field results in a higher failure probability. Figure 0(b) also suggests that negative pairwise interaction provides a buffer for a building against the risk field. This is because to achieve the same mean damage $\mathbb{E}\left[X_{i}\right]>0$, negative $J$ requires larger $h$. In contrast, positive pairwise interaction increases a building’s sensitivity to the risk field, amplifying its tendency towards failure.
@@ 5.1 Example 1: Ising model from post-earthquake damage evaluation data
A neighborhood in Antakya, Turkey, is selected as the target region for constructing the Ising model. The data on building damages is obtained from the Humanitarian OpenStreetMap (HOTOSM) Team [50, 51]. Figure 5 illustrates the damage states for the target region following the 2023 Turkey-Syria earthquake. The region consists of 156 buildings, 40 of which failed during the earthquake.
Figure 5: The building damages in a region of Antakya, Turkey, following the 2023 Turkey-Syria earthquake.
To construct the Ising model, we use the following information: the total number of failed buildings, which is 40 out of 156, and the correlation coefficients between each building pair. Given that we have only a single observation under the earthquake event, we assume spatial ergodicity to obtain the correlation coefficients. Specifically, a correlation coefficient function $\rho(d)$, where $d$ is the distance between two buildings, is fitted to the correlation coefficient values derived by spatially averaging the data. The regression law is found to be $\rho(d)={1}/{\left(1+94.9073d\right)}$, which closely matches the data, as illustrated in Figure 6.
@@
The Ising model is then constructed by Boltzmann machine learning. Figure 7 illustrates the estimated Ising model parameters. Note that we have only one $h$ representing the risk field averaged over the region, rather than a fine-grained $h_{i}$ for each building. This is consistent with the limited prior information—the total number of failed buildings.
Figure 7: The estimated Ising model parameters for Antakya, Turkey, following the 2023 Turkey-Syria earthquake: (a) averaged risk field $h$ and (b) pairwise interaction $\boldsymbol{J}$. Due to the limited prior information, only a single $h$ representing the average risk field is obtained. However, $\boldsymbol{J}$ includes interactions for all pairs of buildings, facilitated by the correlation function $\rho(d)$.
The constructed Ising model reproduces the expected number of failed buildings as 39 (technically, 38.9781), which is close to the observed 40. The validity of the constructed Ising model in terms of reproducing the input information can also be seen in Figure 8. The correlation coefficients reproduced using the Ising model, shown as Figure 7(b), are close to the given values presented in Figure 7(a). It is also noteworthy that the pattern of pairwise interactions closely resembles that of the given correlation coefficients. This suggests that, based on the limited observational data, the correlation coefficients between damages of buildings primarily arise from pairwise interactions between buildings rather than from the risk field.
@@
The Ising model represents a joint probability distribution for all modeled structures over a region, providing a complete statistical description. Using the model, a total of $10^{7}$ samples is generated. The failure probability for each building is then obtained, as illustrated in Figure 9. Figure 9 shows that closely spaced buildings exhibit an overall higher failure probability. This is because the risk field is identical for all buildings, whereas the pairwise interactions are stronger in denser regions. It is important to note that the failure probabilities presented in Figure 9 are predictions of the Ising model based on the current state of knowledge.
@@
Figure 10 illustrates a histogram of the number of failed buildings from the $10^{7}$ samples. It is observed that the mode (most probable number of failed buildings) is around 25, although the expected number is 39. Moreover, the probability of more than 100 buildings will fail is 3.85%, which is not negligible. This is because the distribution of the number of failed buildings has a long right tail. Indeed, it is well known that the Ising model exhibits bimodal for the global mean-field (number of failed buildings in this example) when pairwise interactions are large [28]. Although bimodality is only noticeable in Figure 10, in regions with higher correlation coefficients, such as residential zones with similar houses, the bimodality would be more pronounced and a collective failure would be non-negligible.
@@ 5.2 Example 2: Ising model based on fragility functions
Figure 12 compares the failure probabilities provided by the R2D tool with the risk field values derived from the Ising model, with the color bars min-max normalized. Notably, if we rank the risks of the buildings according to their failure probabilities and compare these rankings to those based on risk field values, the rankings would differ, as illustrated in Figure 13. Specifically, some buildings are considered more risky based on their failure probabilities than their risk field values, while others are viewed as riskier based on the risk field values. This discrepancy can be revealed by their differences in pairwise interactions $\boldsymbol{J}$, as depicted in Figure 14. For instance, building indexed as 108 stands out in Figure 13 for being below and far from the reference line of equal ranking. From Figure 14, it is noted that building #108 also deviates from the rest by having the lowest average interaction with other buildings. This indicates that, in contrast to other buildings which are more likely to fail alongside their close neighborhoods, building #108 exhibits relatively weaker interaction with others. Therefore, contrasting the risk field $\boldsymbol{h}$ with individual failure probabilities reveals the impact of “collective behavior” among structures. This observation is supported by the mean-field approximation in Eq. (16), where the discrepancy between $h_{i}$ and $m_{i}$ reflects the effect of interactions.
@@
Figure 13: Comparison of the normalized risk field and the normalized failure probabilities. The values are min-max normalized. The black solid line represents the reference for equal ranking by risk field and failure probability. Building #108, marked by red, is identified as an “outlier” because it deviates significantly from the reference line of equal ranking.
Figure 14: The average pairwise interaction value for each building. For the $i$-th building, we average its interaction with other buildings to obtain $\sum_{j}J_{ij}/n$. The average interaction for building # 108 is highlighted in red.
Lastly, we examine the relationship between the risk field and earthquake magnitudes. Beyond the existing scenario of magnitude $7.0$, we construct Ising models for earthquake magnitudes of $6.0$, $8.0$, and $9.0$. The results, presented in Figure 15, show that the right/positive mode of $h_{i}$ becomes more dominant as the magnitude increases. The left/negative mode of $h_{i}$ persists across all magnitudes. This is because the mean failure probability is around $0.5$ even at a magnitude of $9.0$, suggesting a significant inertia towards remaining safe, i.e., negative $h_{i}$, among many structures. The relationship between the pairwise interaction and earthquake magnitudes is illustrated by Figure 16. The results suggest that the dispersion of $J_{ij}$ decreases as the magnitude increases, but the spatial mean of $J_{ij}$ across all pairs is not sensitive to the magnitude. This weak dependency between $\boldsymbol{J}$ and magnitude may be attributed to the fact that $\boldsymbol{J}$ primarily reflects the inherent interdependencies among structures.
@@
Figure 16: The relationship between the pairwise interaction $\boldsymbol{J}$ and earthquake magnitude. The left panel shows the histogram of $J_{ij}$ for four earthquake magnitudes, while the right panel illustrates the spatial mean of $J_{ij}$, $\bar{J}=\sum_{ij}J_{ij}/n^{2}$. It is observed that $J_{ij}$ becomes more concentrated with increasing earthquake magnitude, while $\bar{J}$ decreases slowly with magnitude.
@@ 6 Conclusions
This study introduces the long-range Ising model as a joint probability distribution of damage states of structures for regional seismic simulations. Using the principle of maximum entropy, the long-range Ising model is derived under the constraints of the first- and second-order cross-moments of the damage states. The Ising model parameters, $\boldsymbol{h}$ and $\boldsymbol{J}$, are interpreted as a risk field acting at each structure and the pairwise interaction between structures, respectively. Through simple demonstrative examples, the complex interplay between the Ising model parameters and the mean and correlation coefficient of damage states is illustrated. It is found that the mean damage state of a structure is primarily influenced by the risk field acting on it, although the pairwise interactions also play a role in either sharpening or diminishing this effect. Similarly, the correlation coefficient for the damage states between a pair of structures is influenced by all pairwise interactions involving those structures, with the magnitude of the risk field either amplifying or mitigating this effect. The engineering relevance of the Ising model is further illustrated by two simulation test cases. The first example considers the damage states of buildings in Antakya, Turkey, following the 2023 Turkey-Syria earthquake, the second example studies neighborhood in Pacific Heights, San Francisco, under sythetic earthquake scenarios generated by the Regional Resilience Determination (R2D) tool of the NHERI SimCenter. It is found that for both examples the Ising model can reproduce the given information. More importantly, the Ising model offers insights on the collective behavior of structural failures. Specifically, it is found that the discrepancy between the rankings provided by the individual failure probabilities and the risk field values signifies the impact of collective behavior of failing/not-failing together.
@@
    Development of Renormalization Group [28] method for modeling and understanding the multi-scale behavior of regional seismic responses.
@@
    Ising models for functionalities of civil infrastructures.
@@
    Regression models linking ground motion intensity measures to Ising model parameters.

### v2

@@ 4 Parametric study of the Ising model
Using simple proof-of-concept examples, this section investigates the relationships between the Ising model parameters and typical performance quantities, such as failure probability and the correlation coefficient between damage states of structures.
@@ 4.1 Interpretations of the Ising model parameters
    $\boldsymbol{h}$: We interpret $\boldsymbol{h}$ as a risk field acting on each structure within a region. According to Eq. (9), a structure with a positive $h_{i}$ favors the state of failure, i.e., to have a spin of $x_{i}=+1$, with a larger $h_{i}$ signifying a greater tendency toward failure. Conversely, a negative risk field $h_{i}$ indicates a tendency to remain safe. Given that civil structures are engineered to remain reliable in the absence of hazards, they should inherently possess a negative field $-\mathcal{I}$, where $\mathcal{I}>0$ can be interpreted as the inertia toward a safe state. A hazard imposes an external positive field $h_{H}>0$ on each structure, and the resulting risk field is a balance between the hazard and the inertia, i.e., $h_{i}=h_{H}-\mathcal{I}$, where both $h_{H}$ and $\mathcal{I}$ vary across structures. Here, the decomposition of $h_{i}$ into $h_{H}$ and $\mathcal{I}$ is introduced to facilitate a deeper understanding, not for any computational purpose.
@@
    $\boldsymbol{J}$: This represents the pairwise interaction, reflecting the tendency of two structures to be in the same damage state. The term interaction encompasses all direct and indirect effects that determine how the damage state of one structure can affect the state of another, generated by statistical correlations and (potentially) physical contact actions. From Eq. (9), it is observed that a positive $J_{ij}$ favors a synchronization between states, i.e., the corresponding pair of structures have states $(+1,+1)$ or $(-1,-1)$, with a larger $J_{ij}$ increasing this tendency. On the other hand, a negative $J_{ij}$ favors the damage states being different.
The specific choice of terminology, such as risk field and pairwise interaction, can vary, but the fundamental implications of $\boldsymbol{h}$ and $\boldsymbol{J}$ remain invariant. Recall from Section 3 that the construction of the Ising model relies solely on the first- and second-order cross-moments, which might mislead one to argue that the Ising model cannot offer anything beyond the mean and the covariance. However, it is important to acknowledge that the principle of maximum entropy introduces a crucial additional layer of information. Therefore, the Ising model encompasses more than just the mean and the covariance matrix, allowing inferences about the regional seismic behavior that is not apparent from the data.
@@ 4.2 Dependency between the Ising model parameters and the mean and correlation coefficient
Consider two binary random variables $\boldsymbol{X}=[X_{1},X_{2}]^{\intercal}\in\{-1\text{(safe)},1\text{(failure)}\}^{2}$ representing the damage states of two buildings. The risk field and pairwise interaction are given as $h_{1}=h_{2}=h$ and $J_{12}=J$, respectively. With only 2 variables and thus $2^{2}$ system configurations, the exact mean and correlation coefficient for given $h$ and $J$ values can be computed using Eqs. (5)-(9). The mean value $\mathbb{E}\left[X_{i}\right]$ is directly linked to the failure probability $P_{f}$ by $P_{f}=(\mathbb{E}\left[X_{i}\right]+1)/2$. Note that, due to parameter settings of the current demonstration, the failure probability $P_{f}$ for each building is the same.
The mean value for the damage state of a building is affected both by the risk field acting on the site and the pairwise interaction from the other building. Figure 1(a) depicts the mean damage state of a building with varying $h$ and $J$, and Figure 1(b) illustrates the relationship between the mean value and the risk field $h$ under four different $J$ values. The mean damage state of a building becomes larger as the risk field $h$ increases, reflecting that a stronger risk field results in a higher failure probability. Importantly, Figure 1(b) also suggests that: (i) when $h>0$, negative pairwise interaction provides a buffer against failure, i.e., to achieve the same mean damage $\mathbb{E}\left[X_{i}\right]>0$, negative $J$ requires larger $h$; while (ii) when $h<0$, positive pairwise interaction provides a buffer against failure, i.e., to achieve the same $\mathbb{E}\left[X_{i}\right]<0$, positive $J$ requires larger $h$. This observation has a clear interpretation: (i) when the hazard is weak, adding positive correlation into the system increases the likelihood of a collective safe state that helps to resist failure; while (ii) when the hazard is strong, adding positive correlation triggers a collective failure state that is detrimental. In short, coherency can be both a shield and a vulnerability. Incidentally, the change of behaviors with flipping the sign of $h$ is associated with the first-order phase transition in statistical mechanics. We will examine first- and second-order phase transitions within regional seismic responses in upcoming research.
@@ 5.1 Example 1: Ising model from post-earthquake damage evaluation data
A neighborhood in Antakya, Turkey, is selected as the target region for constructing the Ising model. The data on building damage is obtained from the Humanitarian OpenStreetMap (HOTOSM) Team [55, 56]. Figure 5 illustrates the damage states for the target region following the 2023 Turkey-Syria earthquake. The region consists of 156 buildings, 40 of which failed during the earthquake.
Figure 5: The building damage in a region of Antakya, Turkey, following the 2023 Turkey-Syria earthquake. Out of 156 buildings, 40 failed, including 14 out of 72 failing in the inner area and 26 out of 84 failing in the outer area. The inner and outer areas are defined in the right panel; this definition will facilitate the verification of the Ising model predictions.
To construct the Ising model, we use the following information: the total number of failed buildings, which is 40 out of 156, and the correlation coefficients between each building pair. Given that we have only a single observation under the earthquake event, we assume spatial ergodicity to obtain the correlation coefficients. Specifically, a correlation coefficient function $\rho(d)$, where $d$ is the distance between two buildings, is fitted to the correlation coefficient values obtained by spatially averaging the damage data for $29{,}689$ buildings across the entire Antakya. The regression equation is found to be $\rho(d)={1}/{\left(1+94.9073d\right)}$, which closely matches the data, as illustrated in Figure 6. Finally, the correlation coefficient values for the $156$ buildings in the target region are calculated using $\rho(d)$, as depicted in Figure 8(a).
@@
The Ising model is then constructed by Boltzmann machine learning. Figure 7 illustrates the estimated Ising model parameters. Note that we have only one $h$ representing the global risk field for the target region, rather than a fine-grained $h_{i}$ for each building. This is consistent with the limited prior information—the total number of failed buildings.
Figure 7: The estimated Ising model parameters for Antakya, Turkey, following the 2023 Turkey-Syria earthquake: (a) risk field $h$ and (b) pairwise interaction $\boldsymbol{J}$. The prior information only contains the total number of failed buildings instead of individual failure probabilities; thus, a single $h=-0.017$ representing the global risk field for the target region is obtained. On the other hand, $\boldsymbol{J}$ includes interactions for all pairs of buildings, enabled by the correlation function $\rho(d)$.
The constructed Ising model reproduces the expected number of failed buildings as 39, which is close to the observed 40. The validity of the constructed Ising model in terms of reproducing the input information can also be seen in Figure 8. The correlation coefficients reproduced using the Ising model, shown as Figure 8(b), are close to the given values presented in Figure 8(a). Their similarity is quantified by a correlation coefficient computed from flattening each matrix into a vector, and the result is $0.98$. It is also noteworthy that the pattern of pairwise interactions closely resembles that of the given correlation coefficients. This is because, conditional on the limited input information, the correlation coefficients between damage of buildings are primarily driven by pairwise interactions rather than the global risk field.
@@
The Ising model represents a joint probability distribution for all modeled structures over a region, providing a complete statistical description of damage states. Using the model, a total of $10^{7}$ samples is generated. The failure probability for each building is then obtained, as illustrated in Figure 9. Figure 9 shows that closely spaced buildings generally exhibit lower failure probabilities, consistent with the real observation in Figure 5, where about 19% (14 out of 72) of the buildings failed in the inner area while about 31% (26 out of 84) failed in the outer area. This result is consistent with the findings of Figure 1. When the risk field is negative, strong pairwise interactions between closely spaced buildings create a buffer against failure. Conversely, this trend reverses when the risk field is positive.
@@
Figure 10 illustrates a histogram of the number of failed buildings from the $10^{7}$ samples. It is observed that the mode (most probable number of failed buildings) is around 25, although the expected number is 39. Moreover, the probability of more than 100 buildings will fail is 3.85%, which is not negligible. This is because the distribution of the number of failed buildings has a long right tail. Indeed, it is well known that the Ising model exhibits a bimodal distribution for the global mean-field (number of failed buildings in this example) when pairwise interactions are large [33]. Although bimodality is only noticeable in Figure 10, in regions with higher correlation coefficients, such as residential zones with similar houses, the bimodality would be more pronounced and a collective failure would be non-negligible.
@@
The inferences drawn from Figures 9 and 10 exemplify how the Ising model can be used to organize raw data and draw inferences from it. It is important to note that these inferences are conditional on the current state of knowledge; they may not reflect the truth associated with an idealized state of sufficient knowledge. However, inferences from the Ising model can be improved as higher-fidelity information about the target region becomes available.
@@ 5.2 Example 2: Ising model based on fragility functions
Figure 12 compares the failure probabilities provided by the R2D tool with the risk field values derived from the Ising model, with the color bars indicating the rankings of buildings based on these two quantities, where $1$ denotes the highest and $182$ the lowest. The spatial patterns are overall similar, but differences are noticeable. These differences are highlighted in Figure 13 by comparing the risk fields to the failure probabilities, with the reference curve indicating independence between damage states. If all damage states are independent, the points should be distributed along the reference curve and the rankings from the two metrics will be identical. Additionally, if all points follow a monotonically increasing trend, the rankings are identical. In Figure 13, however, the risk field value varies significantly even for similar failure probabilities. This multiplicity is due to the interactions among structures. For instance, the building indexed as $108$ stands out in Figure 13 as being distinctly close to the reference curve, implying that it is closest to independence. In fact, Figure 14 shows that building #108 deviates from the rest by having the smallest average pairwise interaction value. This suggests that the failure event of building #108 has a minimal effect on the state of the other buildings. These observations are supported by the mean-field approximation in Eq. (16), where the discrepancy between $h_{i}$ and $m_{i}$ is contributed by the interaction terms. Therefore, contrasting the risk field $\boldsymbol{h}$ with individual failure probabilities reveals the dominance of collective behavior among structures. The collective behavior is realized by the term ${\boldsymbol{x}}^{\intercal}\boldsymbol{J}\boldsymbol{x}$ in the Ising model, which directly alters the joint distribution; the effect reaches beyond second-order statistics.
@@
Figure 13: Comparison between the risk fields and the failure probabilities. The black solid curve in the left panel represents the exact relationship between $h$ and $P_{f}$ when the damage states are independent, as a special case of the mean-field approximation represented by Eq. (16). The distance from each point to the reference curve of independence is visualized as a solid thin line, which is highlighted in the right panel. Building #108, marked in red, is identified as an outlier because, compared to other buildings with similar failure probabilities, its distance to the reference curve is distinctly small.
Figure 14: The average pairwise interaction value for each building. For the $i$-th building, we average its interaction with other buildings to obtain $\sum_{j\neq i}J_{ij}/\left(n-1\right)$. The average interaction for building #108 is highlighted in red.
Lastly, we examine the relationship between the risk field and earthquake magnitudes. Beyond the existing scenario of magnitude $7.0$, we construct Ising models for earthquake magnitudes of $6.0$, $8.0$, and $9.0$. The results, presented in Figure 15, show that the right/positive mode of $h_{i}$ becomes more dominant as the magnitude increases. The left/negative mode of $h_{i}$ persists across all magnitudes. This is because the mean failure probability is around $0.5$ even at a magnitude of $9.0$, suggesting a significant inertia toward remaining safe, i.e., negative $h_{i}$, among many structures. The relationship between the pairwise interaction and earthquake magnitudes is illustrated by Figure 16. The results suggest that the dispersion of $J_{ij}$ decreases as the magnitude increases, but the spatial mean of $J_{ij}$ across all pairs is not sensitive to the magnitude. This weak dependency between $\boldsymbol{J}$ and magnitude may be attributed to the fact that $\boldsymbol{J}$ primarily reflects the inherent interdependencies among structures.
@@
Figure 16: The relationship between the pairwise interaction $\boldsymbol{J}$ and earthquake magnitude. The left panel shows the histogram of $J_{ij}$ for four earthquake magnitudes, while the right panel illustrates the spatial mean of $J_{ij}$, $\bar{J}=\sum_{i\neq j}J_{ij}/\left(n-1\right)^{2}$. It is observed that $J_{ij}$ becomes more concentrated with increasing earthquake magnitude, while $\bar{J}$ decreases slowly with magnitude.
@@ 6 Conclusions
This study introduces the long-range Ising model as a joint probability distribution of damage states of structures for regional seismic analysis. Using the principle of maximum entropy, the long-range Ising model is derived under the constraints of the first- and second-order cross-moments of the damage states. The Ising model parameters, $\boldsymbol{h}$ and $\boldsymbol{J}$, are interpreted as a risk field acting at each structure and the pairwise interaction between structures, respectively. Through simple demonstrative examples, the complex interplay between the Ising model parameters and the mean and correlation coefficient of damage states is illustrated. It is found that the mean damage state of a structure is primarily influenced by the risk field acting on it, although the pairwise interactions also play a role in either sharpening or diminishing this effect. Similarly, the correlation coefficient for the damage states between a pair of structures is influenced by all pairwise interactions involving those structures, with the magnitude of the risk field either amplifying or mitigating this effect. The engineering relevance of the Ising model is further illustrated by two simulation test cases. The first example considers the damage states of buildings in Antakya, Turkey, following the 2023 Turkey-Syria earthquake, the second example studies neighborhood in Pacific Heights, San Francisco, under sythetic earthquake scenarios generated by the Regional Resilience Determination (R2D) tool of the NHERI SimCenter. It is found that for both examples the Ising model can reproduce the given information. More importantly, the Ising model offers insights on the collective behavior of structural failures. Specifically, it is found that the discrepancy between the risk field values and the individual failure probabilities signifies the dominance of collective behavior in group failures or non-failures.
@@
    Development of Renormalization Group [33] method for modeling and understanding the multi-scale behavior and phase transitions in regional seismic responses.
@@
    Ising models for multiple damage states and functionalities of networked civil infrastructures.
@@
    Comprehensive comparative studies using high-fidelity data on regional seismic responses.